OpenStudy (anonymous):

Solve the equation for x: 3x + y = 6. Show your work.

OpenStudy (anonymous):

3x=-y+6 -- -- -- 3 3 3 x=-y/3+2

OpenStudy (anonymous):

3x+y=6 3x=6-y x=(6-y)/3